# WorldSense: Evaluating Real-world Omnimodal Understanding for Multimodal LLMs

### v1

@@ 3.5 Evaluation Paradigm
In our evaluation framework, each test instance consists of a video clip with synchronized audio and a multiple-choice question. Models must process these multi-modal inputs and select the correct answer from several options. Performance is measured by accuracy, comparing the model’s selection to the ground-truth answers. A model’s success is determined by its ability to accurately align with the correct answer.
@@ 4 Experiments and Findings
This section presents a comprehensive evaluation of both open-source and proprietary MLLMs on the WorldSense benchmark. We first delineate our experimental methodology and evaluation protocols, followed by a comprehensive analysis of quantitative results. Furthermore, we conduct detailed investigations into the important factors that affect performance, providing insights that illuminate potential directions for multi-modal understanding.
@@ 4.2 Results on WorldSense
Main Results. We present the comprehensive evaluation results on WorldSense in Table 2. Our findings reveal several significant insights regarding the current state of multi-modal models in real-world understanding.
@@
Methods Modality Tech & Science Culture & Politics Daily Life Film & TV Performance Games Sports Music Avg Unified-IO-2 L [42] Audio 23.0 25.4 24.2 26.7 27.7 23.7 25.0 27.1 25.2 + Caption 21.5 21.1 20.7 17.1 19.9 19.0 22.9 23.7 20.9-4.3 + Video 19.3 22.8 23.1 25.6 25.8 24.1 22.9 25.3 23.3-1.9 Unified-IO-2 XL [42] Audio 21.7 22.4 22.4 22.1 24.7 25.0 25.9 24.7 23.4 + Caption 19.9 19.8 20.8 19.2 20.2 15.9 21.7 25.5 20.7-2.7 + Video 26.5 24.4 22.5 23.5 24.7 28.0 25.7 24.2 24.7+1.3 Unified-IO-2 XXL [42] Audio 27.5 28.7 23.9 23.2 25.8 21.1 26.2 30.2 25.9 + Caption 24.0 26.7 23.0 18.9 18.7 20.7 25.9 29.4 23.7-2.2 + Video 27.1 31.7 23.9 23.7 25.5 23.7 25.7 27.3 25.9+0.0 OneLLM [21] Audio 25.7 26.1 19.3 21.9 25.8 25.9 21.5 22.4 23.0 + Caption 29.6 29.0 25.9 29.1 33.0 26.7 29.2 28.6 28.6+5.6 + Video 26.7 25.1 19.0 22.7 27.0 23.7 22.4 19.8 22.8-0.2 VideoLLaMA2 [7] Audio 23.8 23.4 21.3 22.4 24.7 19.8 27.1 27.9 23.8 + Caption 30.0 30.0 25.6 29.9 28.5 25.0 29.7 29.9 28.5+4.7 + Video 29.4 25.4 21.8 24.5 26.2 24.6 25.5 27.1 25.4+1.6 Gemini 1.5 Pro [58] Audio 40.2 42.9 35.8 33.3 33.0 31.0 33.3 24.7 34.6 + Caption 49.5 52.1 41.8 42.9 46.4 41.8 39.6 36.7 43.6+9.0 + Video 53.7 47.2 50.3 50.4 52.4 46.8 40.2 42.0 48.0+13.4
Second and surprisingly, existing open-source audio-visual MLLMs perform even worse, achieving accuracy rates comparable to random guessing and notably below video-only MLLMs. This counter-intuitive finding reveals that despite having access to both modalities, these models struggle with effective audio-visual integration, suggesting that multimodal processing capability alone does not guarantee better performance without sophisticated integration mechanisms.
Third, among proprietary MLLMs, vision-only models GPT-4o and Claude 3.5 Sonnet perform similarly to the best open-source video MLLMs. Gemini 1.5 Pro, capable of processing both audio and visual information, achieves the highest accuracy of . However, this performance still falls considerably short of requirements for reliable real-world applications, indicating substantial room for improvement.
@@
Breakdown Results. We conduct a fine-grained analysis of model performance across different audio types and task categories, as shown in Figure 4, revealing important insights into the limitations of existing multimodal models.
@@
Additionally, performance varies across audio types. While Gemini 1.5 Pro performs best overall, it shows notably lower accuracy on event-related questions compared to speech or music tasks, possibly due to the complex nature of environmental sounds. Other models also exhibit inconsistent performance across audio types, underscoring a general limitation for existing models for audio understanding.
@@ 4.3 Roadmap Towards Real-world Understanding
The results in Table 5 reveal intriguing patterns in how different forms of audio information influence model performance. For Gemini 1.5 Pro, accuracy increases from (video-only) to with subtitles, and further to with original audio. OneLLM shows similar improvements. These results demonstrate that both subtitles and acoustic features (including tone, emotion, and environmental sounds) contribute valuable information for multimodal understanding, beyond what subtitles alone can capture, highlighting the importance of complete acoustic cues for multimodal understanding.
Interestingly, UnifiedIO2 demonstrates performance degradation when integrating either subtitles or audio, with subtitles causing a notable accuracy decline, suggesting difficulties in multimodal processing. Conversely, Video-LLaMA2 improves with both modalities but performs better with subtitles than original audio, indicating a stronger reliance on textual rather than complex acoustic information. We further evaluate video-only MLLMs by providing transcribed subtitles, as shown in Table 5. Nearly all models show significant improvements with subtitle integration, reinforcing the importance of audio information. However, the performance gain is less pronounced in music-related questions, as subtitles cannot effectively capture inherent acoustic features such as melody, rhythm, and harmony.
@@
Methods Speech Event Music Overall Video + Subtitle + Audio Video + Subtitle + Audio Video + Subtitle + Audio Video + Subtitle + Audio Unified-IO-2 L [42] 26.8 13.9-12.9 23.1-3.1 26.9 13.5-13.4 23.2-3.7 26.3 15.0-11.3 24.5-1.8 26.6 14.8-11.8 23.3-3.3 Unified-IO-2 XL [42] 25.0 13.0-12.0 24.6-0.4 24.8 12.3-12.5 25.0+0.2 26.7 15.9-10.8 24.7-2.0 25.3 14.1-11.2 24.7-0.6 Unified-IO-2 XXL [42] 27.0 15.6-11.4 25.7-1.3 26.2 14.2-12.0 25.1-1.1 28.4 19.1-9.3 25.9-2.5 27.2 17.2-10.0 25.9-1.3 OneLLM [21] 12.5 19.6+7.1 23.2+10.7 12.4 19.3+6.9 22.3+9.9 12.4 19.0+6.6 22.5+10.1 12.6 19.6+7.0 22.8+10.2 VideoLLaMA2 [7] 17.1 25.5+8.4 25.2+8.1 16.1 24.9+8.8 24.5+8.4 17.7 27.0+9.3 25.6+7.9 17.4 26.1+8.7 25.4+8.0 Gemini 1.5 Pro [58] 34.3 39.6+5.3 49.2+14.9 33.0 38.9+5.9 48.7+15.7 35.4 39.2+3.8 49.8+14.4 34.4 39.3+4.9 48.0+13.6
@@
Video Frames. We investigate the effect of temporal sampling density by varying input frame numbers for video-only MLLMs. As shown in Figure 5, most models demonstrate significant performance improvements with increased frame density, with LLaMA-3.2 being a notable exception. These gains likely result from better capture of fine-grained temporal dynamics and subtle visual changes, emphasizing the importance of dense temporal sampling.
@@ 5 Conclusion
In this paper, we introduce WorldSense, the first benchmark designed to evaluate MLLMs’ omnimodal understanding in real-world scenarios. Distinguished by its emphasis on joint omnimodal comprehension across diverse real-world contexts, WorldSense encompasses rich video categories and carefully curated question-answer pairs that necessitate the integration of visual and acoustic information. Through extensive experiments, we expose significant limitations in current MLLMs’ ability to process and coherently integrate omnimodal information. Our analysis demonstrates the importance of omnimodal collaboration in real-world understanding. We hope that WorldSense can serve as a foundational benchmark for advancing human-like omnimodal understanding capabilities.
Appendix A Broader Impacts
@@
Our work on WorldSense has several potential positive impacts on society and AI development, while also presenting certain risks that warrant careful consideration. WorldSense contributes to advancing MLLMs’ ability to understand and interact with the real world through multiple modalities. This progress could benefit various applications, including assistive technologies, educational tools, human-AI interaction systems, safety systems, and so on. We also acknowledge potential risks and challenges. The development of more capable AI systems might raise privacy concerns. Advanced multimodal understanding capabilities could potentially be misused for surveillance or monitoring purposes. We believe that open discussion of these impacts is crucial for responsible development of MLLMs.
Appendix B Ethics Statement
Our research on WorldSense adheres to strict ethical principles and guidelines. We acknowledge several important ethical considerations:
Data Collection and Privacy. All video content in WorldSense has been collected from publicly available sources with appropriate licensing agreements. We have con ducted thorough reviews and implemented comprehensive data processing procedures to ensure privacy protection, including the removal of any personally identifiable information.
Potential Biases. While acknowledging that inherent biases may exist in any dataset, we have undertaken systematic efforts to ensure diverse representation across our video content and question-answer pairs, encompassing various domains, cultures, and contexts. Nevertheless, we recognize that completely eliminating bias remains a significant challenge, and users should carefully consider these potential limitations when utilizing our dataset.
Intended Use. WorldSense is specifically designed to advance research in omnimodal real-world understanding. While we actively encourage the use of this benchmark for academic and research purposes, we strongly caution against any applications that could potentially result in harmful or discriminatory outcomes. Users are expected to adhere to ethical guidelines and responsible practices.
Appendix C License

### v2

@@ 3.5 Evaluation Paradigm
In our evaluation framework, each test instance consists of a video clip with synchronized audio and a multiple-choice question. Models must process these multi-modal inputs and select the correct answer from several options. Performance is measured by accuracy, comparing the model’s selection to the ground-truth answers. A model’s success is determined by its ability to accurately align with the correct answer. We employ a matching-based approach to extract answers.
@@ 4.1 Settings
Methods LLM Size Tech & Science Culture & Politics Daily Life Film & TV Perfor- mance Games Sports Music Avg Open-Source Video-Audio MLLMs Unified-IO-2 L [43] 1B 19.3 22.8 23.1 25.6 25.8 24.1 22.9 25.3 23.3 Unified-IO-2 XL [43] 3B 26.5 24.4 22.5 23.5 24.7 28.0 25.7 24.2 24.7 Unified-IO-2 XXL [43] 7B 27.1 31.7 23.9 23.7 25.5 23.7 25.7 27.3 25.9 OneLLM [22] 7B 26.7 25.1 19.0 22.7 27.0 23.7 22.4 19.8 22.8 VideoLLaMA2 [7] 7B 29.4 25.4 21.8 24.5 26.2 24.6 25.5 27.1 25.4 VITA-1.5 [18] 7B 38.2 35.9 34.3 39.8 41.2 32.6 34.7 39.9 36.9 Qwen2.5-Omni [71] 7B 47.8 49.8 43.6 43.8 48.3 39.1 43.5 47.3 45.4 Open-Source Video MLLMs Video-LLaVA [33] 7B 23.6 20.8 19.1 17.3 23.6 17.2 20.8 20.1 20.3 LLaMA3.2 [21] 7B 27.5 25.7 28.9 25.9 27.7 21.1 29.0 26.8 27.1 Qwen2-VL [64] 7B 33.5 29.0 28.4 33.6 30.3 32.3 34.7 38.5 32.4 mPLUG-Owl3 [75] 7B 37.5 31.4 31.0 34.1 33.3 33.2 32.1 30.5 32.9 LLaVA-OneVision [27] 7B 38.9 38.9 36.3 37.6 37.8 37.9 36.3 39.1 37.7 InternVL2.5 [5] 8B 43.7 40.9 34.6 39.7 37.8 36.2 39.4 41.1 39.1 LLaVA-Video [82] 7B 41.6 38.6 40.6 42.1 40.4 39.7 37.0 40.9 40.2 Proprietary MLLMs Claude 3.5 Sonnet [2] - 43.7 31.7 30.6 36.5 30.7 31.9 36.6 33.9 34.8 GPT 4o [25] - 48.0 44.0 38.3 43.5 41.9 41.2 42.6 42.7 42.6 Gemini 1.5 Pro [60] - 53.7 47.2 50.3 50.4 52.4 46.8 40.2 42.0 48.0
@@ 4.2 Results on WorldSense
Main Results. We present comprehensive evaluations of WorldSense in Table 2. Our analysis reveals several significant insights regarding the current capabilities of MLLMs in real-world understanding.
@@
Second and surprisingly, most of existing open-source audio-visual MLLMs perform even worse, achieving accuracy rates comparable to random guessing and notably below video-only MLLMs. This counter-intuitive finding reveals that despite having access to both modalities, these models struggle with effective audio-visual integration, suggesting that multimodal processing capability alone does not guarantee better performance without sophisticated integration mechanisms.
Third, among proprietary MLLMs, vision-only models GPT-4o and Claude 3.5 Sonnet demonstrate performance comparable to the leading open-source video MLLMs. Gemini 1.5 Pro, capable of processing both audio and visual information, achieves the highest accuracy of . However, this performance still falls considerably short of requirements for reliable real-world applications, indicating substantial room for improvement.
@@
Breakdown Results. We conduct a fine-grained analysis of model performance across different audio types and task categories, as shown in Figure 4 and 7, highlighting the limitations of MLLMs.
@@
Additionally, performance varies across audio types. While Gemini 1.5 Pro performs best overall, it shows notably lower accuracy on event-related questions compared to speech or music tasks, possibly due to the complex nature of environmental sounds. Other models also exhibit inconsistent performance across audio types, underscoring a general limitation for audio understanding.
Methods Modality Tech & Science Culture & Politics Daily Life Film & TV Perfor- mance Games Sports Music Avg Unified-IO-2 L [43] Audio 23.0 25.4 24.2 26.7 27.7 23.7 25.0 27.1 25.2 + Caption 21.5 21.1 20.7 17.1 19.9 19.0 22.9 23.7 20.9-4.3 + Video 19.3 22.8 23.1 25.6 25.8 24.1 22.9 25.3 23.3-1.9 Unified-IO-2 XL [43] Audio 21.7 22.4 22.4 22.1 24.7 25.0 25.9 24.7 23.4 + Caption 19.9 19.8 20.8 19.2 20.2 15.9 21.7 25.5 20.7-2.7 + Video 26.5 24.4 22.5 23.5 24.7 28.0 25.7 24.2 24.7+1.3 Unified-IO-2 XXL [43] Audio 27.5 28.7 23.9 23.2 25.8 21.1 26.2 30.2 25.9 + Caption 24.0 26.7 23.0 18.9 18.7 20.7 25.9 29.4 23.7-2.2 + Video 27.1 31.7 23.9 23.7 25.5 23.7 25.7 27.3 25.9+0.0 OneLLM [22] Audio 25.7 26.1 19.3 21.9 25.8 25.9 21.5 22.4 23.0 + Caption 29.6 29.0 25.9 29.1 33.0 26.7 29.2 28.6 28.6+5.6 + Video 26.7 25.1 19.0 22.7 27.0 23.7 22.4 19.8 22.8-0.2 VideoLLaMA2 [7] Audio 23.8 23.4 21.3 22.4 24.7 19.8 27.1 27.9 23.8 + Caption 30.0 30.0 25.6 29.9 28.5 25.0 29.7 29.9 28.5+4.7 + Video 29.4 25.4 21.8 24.5 26.2 24.6 25.5 27.1 25.4+1.6 VITA-1.5 [18] Audio 30.2 35.6 36.3 30.9 32.2 32.2 31.4 33.3 32.9 + Caption 39.2 39.8 37.2 37.5 37.5 35.2 34.9 38.4 37.5+4.6 + Video 38.2 35.9 34.3 39.8 41.2 32.6 34.7 39.9 36.9+4.0 Qwen2.5-Omni [71] Audio 40.0 38.2 36.0 33.5 31.1 30.5 32.3 33.3 34.9 + Caption 40.0 37.9 38.9 33.5 36.7 37.8 37.7 38.9 37.9+3.0 + Video 47.8 49.8 43.6 43.8 48.3 39.1 43.5 47.3 45.4+10.5 Gemini 1.5 Pro [60] Audio 40.2 42.9 35.8 33.3 33.0 31.0 33.3 24.7 34.6 + Caption 49.5 52.1 41.8 42.9 46.4 41.8 39.6 36.7 43.6+9.0 + Video 53.7 47.2 50.3 50.4 52.4 46.8 40.2 42.0 48.0+13.4
@@ 4.3 Roadmap Towards Real-world Understanding
The results in Table 4 reveal intriguing patterns in how different forms of audio information influence model performance. For Gemini 1.5 Pro, accuracy increases from (video-only) to with subtitles, and further to with original audio. Other models, such as OneLLM and Qwen2.5-Omni, show similar improvements. These results demonstrate that both subtitles and acoustic features (including tone, emotion, and environmental sounds) contribute valuable information, beyond what subtitles alone can capture, emphasizing the importance of complete acoustic cues in omni-modal real-world understanding.
Interestingly, UnifiedIO2 demonstrates performance degradation when integrating either subtitles or audio, with subtitles causing a notable accuracy decline, suggesting difficulties in multimodal processing. Conversely, Video-LLaMA2 improves with both modalities but performs better with subtitles than original audio, indicating a stronger reliance on textual rather than acoustic information.
We further evaluate video-only MLLMs by providing transcribed subtitles, as shown in Table 4. Nearly all models show significant improvements with subtitle integration, reinforcing the importance of audio information. However, the performance gain is less pronounced in music-related questions, as subtitles cannot effectively capture inherent acoustic features such as melody, rhythm, and harmony.
@@
Methods Speech Event Music Overall Video + Subtitle + Audio Video + Subtitle + Audio Video + Subtitle + Audio Video + Subtitle + Audio Unified-IO-2 L [43] 26.8 13.9-12.9 23.1-3.1 26.9 13.5-13.4 23.2-3.7 26.3 15.0-11.3 24.5-1.8 26.6 14.8-11.8 23.3-3.3 Unified-IO-2 XL [43] 25.0 13.0-12.0 24.6-0.4 24.8 12.3-12.5 25.0+0.2 26.7 15.9-10.8 24.7-2.0 25.3 14.1-11.2 24.7-0.6 Unified-IO-2 XXL [43] 27.0 15.6-11.4 25.7-1.3 26.2 14.2-12.0 25.1-1.1 28.4 19.1-9.3 25.9-2.5 27.2 17.2-10.0 25.9-1.3 OneLLM [22] 12.5 19.6+7.1 23.2+10.7 12.4 19.3+6.9 22.3+9.9 12.4 19.0+6.6 22.5+10.1 12.6 19.6+7.0 22.8+10.2 VideoLLaMA2 [7] 17.1 25.5+8.4 25.2+8.1 16.1 24.9+8.8 24.5+8.4 17.7 27.0+9.3 25.6+7.9 17.4 26.1+8.7 25.4+8.0 VITA-1.5 [18] 37.6 39.1+1.5 36.2-1.4 36.4 38.2+1.8 35.3-1.1 38.7 40.0+1.3 37.1-1.6 37.7 39.3+1.6 36.5-1.2 Qwen2.5-Omni [71] 38.7 38.7+0.0 44.8+6.1 37.6 37.7+0.1 43.8+6.2 40.7 40.3-0.4 46.1+5.4 39.2 39.2+0.0 45.2+6.0 Gemini 1.5 Pro [60] 34.3 39.6+5.3 49.2+14.9 33.0 38.9+5.9 48.7+15.7 35.4 39.2+3.8 49.8+14.4 34.4 39.3+4.9 48.0+13.6
@@
Failure Analysis and Future Improvement. We perform error analysis on 130 samples of Gemini 1.5 Pro (5 random samples per task) through manual review, identifying three main error types: Audio Understanding Errors (misinterpreting audio information), Visual Understanding Errors (missing visual details), and Reasoning Errors (faulty logical steps). As shown in Figure 7, most errors stem from audio understanding deficiencies and reasoning failures. The reason for poor accuracy and limitation of existing models can be summarized as follows: (i) Inadequate Audio Understanding. Existing models fail to understand audio information correctly and show significantly weaker audio processing than visual understanding. (ii) Limited Cross-Modal Integration. Models often process modalities independently rather than performing true multimodal integration and suffer from insufficient omni-modal information integration. (iii) Insufficient Complex Reasoning Ability. Despite correct perception, MLLMs still conduct error reasoning, leading to incorrect conclusions.
We also raise several key strategies to enhance models’ understanding of omni-modality information: (i) Coupled Multimodal Training Data. Using naturally coupled, interleaved multimodal data, for example, audio, visual, language content, would enhance models’ capability to leverage cross-modal dependencies. (ii) Architectural Improvements. Enhanced attention mechanisms facilitating deep multimodal integration could emphasize early fusion between modalities, rather than processing them as separate streams for late fusion. (iii) Advanced Modal Alignment Techniques. Progressive alignment strategies that gradually enhance the model’s ability to align information across modalities could lead to more effective utilization of multimodal inputs. (iiii) Reasoning strengthening. Incorporating diverse reasoning-focused data can strengthen logical inference capabilities, enabling more coherent and accurate conclusions.
@@ 5 Conclusion
In this paper, we propose WorldSense, the first benchmark designed to evaluate MLLMs’ omnimodal understanding in real-world scenarios. Distinguished by its emphasis on joint omnimodal comprehension across diverse real-world contexts, WorldSense encompasses rich video categories and carefully curated question-answer pairs that necessitate the integration of visual and acoustic information. Through extensive experiments, we expose significant limitations in current MLLMs’ ability to process and coherently integrate omnimodal information. Our analysis demonstrates the importance of omnimodal collaboration in real-world understanding. We hope that WorldSense can serve as a foundational benchmark for advancing human-like omnimodal understanding capabilities.
Appendix A Quality Control
Experienced annotators. Our annotation team consists of 80 professional annotators with extensive QA annotation experience. These annotators are proficient in English, and have participated in sereval QA data annotation projects.
Sufficient annotation training. We conducted a one-week training program with 200 videos (excluded from final benchmark) until annotators achieved high proficiency (only requiring modifications).
Annotation instruction. Each annotator received a comprehensive instruction with task explanations, question formulation guidelines, QA creation instructions, annotated examples, and cross-modal inference requirements.
Our review process identified and revised similar QA pairs. Through professional annotators, thorough training, detailed guidelines, and rigorous quality control, we ensure high-quality annotations.
Appendix B Implement Details
For open-source MLLMs, we strictly follow their official implementations and recommended preprocessing pipelines to ensure fair comparison. For GPT 4o and Claude 3.5 Sonnet, we sample 16 frames uniformly from each video, while for Gemini 1.5 Pro, we utilize the official API for raw video file uploads. We conduct all the experiments on a NVIDIA A100 GPU.
Appendix C Evaluation Prompt
Following previous works [17, 31], we adopt the format of “whole video frames + whole subtitles/audios (optional) + question with prompt” as prompt. We show the evaluation prompt across three input configurations: video-only input, video with subtitles, and video with audio content as following.
Prompt Avg Prompt 1 24.0 Prompt 2 24.8 Prompt Ours 25.4
Appendix D Prompt Sensitivity
We have conducted on experiments on VideoLLaMA2 to explore the impact of different prompts, and results are shown in Table 5. Prompt 1 stands for "Based on your observations, select the best option that accurately addresses the question.", Prompt 2 is "Carefully watch this video and pay attention to every detail.", and Prompt Ours represents the prompt in Section C. As shown in the experiment, the prompt template we ultimately used consistently yielded superior performance.
Appendix E Failure Case
Figure 8 showcases two failure cases of Gemini 1.5 Pro, VideoLLaMA2, and LLaVA-Video. The left case involves a vision understanding and reasoning error, where Gemini 1.5 Pro incorrectly identified the clock time as 10:30 instead of the actual 10:00 displayed, leading to the incorrect answer. This reflects deficiencies in basic visual perception and properly correlating visual information with the question. The right case demonstrates an audio understanding error, where models misjudged the rhythm pattern change in guzheng music, interpreting it as changing from soothing to intense (option D) rather than the correct passionate to soothing (option B). This case indicates that tasks involving interpretation of musical emotion and rhythm patterns remain challenging for existing MLLMs. These cases highlight the current limitations of MLLMs in processing subtle visual details and interpreting subjective audio qualities.
Appendix F Limitation
While our WorldSense represents a significant advancement in evaluating multimodal understanding capabilities of MLLMs, the multiple-choice format inevitably constrains the assessment of models’ generative capabilities. Real-world understanding often requires open-ended responses, explanations, and adaptability beyond selecting from predefined options. Our WorldSense may not adequately evaluate how models perform on tasks requiring nuanced reasoning or creative problem-solving. We will add open-ended questions and expand the evaluation paradigm to better assess real-world multimodal understanding.
Appendix G Broader Impacts & Ethics Statement
@@
Our work on WorldSense has several potential positive impacts on society and AI development, while also presenting certain risks that warrant careful consideration. WorldSense contributes to advancing MLLMs’ ability to understand and interact with the real world through multiple modalities. This progress could benefit various applications, including assistive technologies, educational tools, human-AI interaction systems, safety systems, and so on. We also acknowledge potential risks and challenges. The development of more capable AI systems might raise privacy concerns. Advanced multimodal understanding capabilities could potentially be misused for surveillance or monitoring purposes. We believe that open discussion of these impacts is crucial for the responsible development of multi-modal large language models.
Our research on WorldSense adheres to strict ethical principles and guidelines. We acknowledge several important ethical considerations: (1) Data Collection and Privacy. All video content in WorldSense has been collected from publicly available sources with appropriate licensing agreements. We have con ducted thorough reviews and implemented comprehensive data processing procedures to ensure privacy protection, including the removal of any personally identifiable information. (2) Potential Biases. While acknowledging that inherent biases may exist in any dataset, we have undertaken systematic efforts to ensure diverse representation across our video content and question-answer pairs, encompassing various domains, cultures, and contexts. Nevertheless, we recognize that completely eliminating bias remains a significant challenge, and users should carefully consider these potential limitations when utilizing our dataset. (3) Intended Use. WorldSense is specifically designed to advance research in omnimodal real-world understanding. While we actively encourage the use of this benchmark for academic and research purposes, we strongly caution against any applications that could potentially result in harmful or discriminatory outcomes. Users are expected to adhere to ethical guidelines and responsible practices.
Appendix H License